In a survey of 500 infants chosen at random, it is found that 240 are girls. Are boy and girl births equally likely according to this survey (use α = 0.05)


1
Expert's answer
2021-07-15T10:09:44-0400

H0: there is no significant difference between boy births and girl births

H1: there is a significant difference between boy births and girl births

The probability of girl birth "= p =\frac{1}{2}=0.5"

The probability of boy birth "= q = \frac{1}{2}=0.5"

"n=500 \\

\sqrt{ \frac{pq}{n} }= \sqrt{ \frac{0.5 \times 0.5}{500} }=0.02236"

The observed sample proportion of girl birth

"\hat{p}= \frac{240}{500}=0.48"

The test-statistic:

"Z = \frac{\hat{p}-p}{\sqrt{ \frac{pq}{n} }} \\

Z = \frac{0.48-0.5}{0.02236}= -0.894 \\

α=0.05 \\

Z_{crit}= 1.96 \\

Z < Z_{crit}"

Accept H0.

There is enough evidence to conclude that here is no significant difference between boy births and girl births at the 0.05 significance level.